# An Integral Equation Arising from Time-Consistent Portfolio Selection

### v1

@@ 4.1 Dynamic Mean-Variance portfolio selection
In this subsection, we examine the problem that was studied by Björk, Murgoci, and Zhou (2014), [^5] The trading strategies in Björk, Murgoci, and Zhou (2014) is defined as the dollar amounts put into the risky assets. Their equilibrium strategy is formally a proportion strategy, but any perturbation of this strategy, which is a dollar-amount strategy, can lead to negative wealth. Therefore, as He and Jiang (2021, Page 3872) pointed out, the definition of the equilibrium strategy in Björk, Murgoci, and Zhou (2014) needs a modification such that their solution is indeed an equilibrium strategy. Our model, however, uses the proportion strategies to avoid this issue. where the MV portfolio selection was investigated under the assumption that the agent’s risk aversion is inversely proportional to the wealth level. Specifically, the agent’s MV preference functional at time $t$ can be equivalently given by (3.3). It is easy to see that $g$ and $h$ are independent of $t$ and, for all $(x,y)\in[0,\infty)\times\mathbb{R}$,
@@
$-v_{a}(t)-y_{a}(t)$ $\leq-v_{a}(t)+\sqrt{v_{\lambda}(0)}\sqrt{v_{a}(t)}$
$=\sqrt{v_{a}(t)}(\sqrt{v_{\lambda}(0)}-\sqrt{v_{a}(t)})\leq\frac{{v_{\lambda}(0)}}{4}.$
@@
$\tilde{a}(t)=\mathcal{P}_{t}\left(\tilde{h}(v_{\tilde{a}}(t),y_{\tilde{a}}(t))\right)=\mathcal{P}_{t}\left({h}(v_{\tilde{a}}(t),y_{\tilde{a}}(t))\right),\quad t\in[0,T).$
@@ 4.2 Equilibrium investment with random risk aversion
$h(x)=\frac{\mathbb{E}\left[\mathrm{e}^{-\frac{1}{2}\bm{R}x^{2}}\right]}{\mathbb{E}\left[\bm{R}\mathrm{e}^{-\frac{1}{2}\bm{R}x^{2}}\right]},\quad x\in[0,\infty).$ (4.3)
To study the Lipschitz continuity of $h$, we introduce the following function:
$f(x)=\mathbb{E}\left[\mathrm{e}^{-\frac{1}{2}\bm{R}x^{2}}\right],\quad x\in[0,\infty).$
The next lemma summarizes some well-known properties of $f$ and $h$; see, Klenke (2020, Example 6.29).
@@
If $\mathbb{E}\left[\bm{R}^{2}\right]<\infty$ and $h$ is bounded, then equilibrium exists uniquely in $\mathcal{D}$.
@@
In the case when $\Gamma$ is a two-point distribution, Desmettre and Steffensen (2023) characterize the equilibrium by a three-dimensional ODE system666 In fact, with appropriate transformations, their ODE system can be converted into our integral equation. without establishing the existence and uniqueness of the solution. Our approach, however, only requires that $\mathbb{E}\left[\bm{R}^{2}\right]<\infty$ and that $h$ is bounded to establish the existence and uniqueness of the solution. The boundedness of $h$ is fulfilled when the support of $\Gamma$ has a positive lower bound, that is, $\Gamma(\gamma_{0})=0$ for some $\gamma_{0}>0$. Empirical studies, in fact, indicate that the risk aversion coefficient is bounded away from 00, as discussed in Chetty (2006) and Gandelman and Hernández-Murillo (2015), and the references therein.
Theorem 4.7.
Suppose that $d=1$, $\lambda(\cdot)>0$ is constant, and there are no portfolio bounds. Then the solution $a$ of (2.1) decreases with $t$. Furthermore, if $\Gamma$ is not a single-point distribution, $a$ is strictly decreasing with $t$.
Proof.
In this case, we know that
$a(t)=\frac{\mathbb{E}\left[\mathrm{e}^{-\frac{1}{2}\bm{R}v(t)}\right]}{\mathbb{E}\left[\bm{R}\mathrm{e}^{-\frac{1}{2}\bm{R}v(t)}\right]}\lambda.$
Then
$a’(t)=-\frac{1}{2}v^{\prime}(t)\frac{\left(\mathbb{E}\left[\bm{R}\mathrm{e}^{-\frac{1}{2}\bm{R}v(t)}\right]\right)^{2}-\mathbb{E}\left[\mathrm{e}^{-\frac{1}{2}\bm{R}v(t)}\right]\mathbb{E}\left[\bm{R}^{2}\mathrm{e}^{-\frac{1}{2}\bm{R}v(t)}\right]}{\left(\mathbb{E}\left[\bm{R}\mathrm{e}^{-\frac{1}{2}\bm{R}v(t)}\right]\right)^{2}}\lambda.$
Note that
$\left(\mathbb{E}\left[\bm{R}\mathrm{e}^{-\frac{1}{2}\bm{R}v(t)}\right]\right)^{2}$ $=\left(\mathbb{E}\left[\sqrt{\mathrm{e}^{-\frac{1}{2}\bm{R}v(t)}}\sqrt{\bm{R}^{2}\mathrm{e}^{-\frac{1}{2}\bm{R}v(t)}}\right]\right)^{2}$
$\leq\mathbb{E}\left[\mathrm{e}^{-\frac{1}{2}\bm{R}v(t)}\right]\mathbb{E}\left[\bm{R}^{2}\mathrm{e}^{-\frac{1}{2}\bm{R}v(t)}\right],$
[…]
$\mathbb{E}\left[\mathrm{e}^{-\frac{1}{2}\bm{R}v(t)}\right]\bm{R}^{2}\mathrm{e}^{-\frac{1}{2}\bm{R}v(t)}=\mathrm{e}^{-\frac{1}{2}\bm{R}v(t)}\mathbb{E}\left[\bm{R}^{2}\mathrm{e}^{-\frac{1}{2}\bm{R}v(t)}\right],$
which is equivalent to $\Gamma$ being a single-point distribution. Therefore, $a$ is strictly decreasing with $t$ unless $\Gamma$ is a single-point distribution. ∎
Remark 4.8.
Suppose that $d=1$, $\lambda(\cdot)>0$ and $\sigma(\cdot)>0$ are constants, and there are no portfolio bounds. Then the equilibrium strategy $\bar{\pi}=\frac{1}{\sigma}a$ decreases with $t$. This indicates that the equilibrium strategy exhibits a horizon effect, which confirms the numerical results in Desmettre and Steffensen (2023).

### v2

@@ 4.1 Dynamic Mean-Variance portfolio selection
In this subsection, we examine the problem that was studied by Björk, Murgoci, and Zhou (2014), [^5] The trading strategies in Björk, Murgoci, and Zhou (2014) is defined as the dollar amounts put into the risky assets. Their equilibrium strategy is formally a proportion strategy, but any perturbation of this strategy, which is a dollar-amount strategy, can lead to negative wealth. Therefore, as He and Jiang (2021, Page 3872) pointed out, the definition of the equilibrium strategy in Björk, Murgoci, and Zhou (2014) needs a modification such that their solution is indeed an equilibrium strategy. Our model, however, uses the proportion strategies to avoid this issue. where the MV portfolio selection was investigated under the assumption that the agent’s risk aversion is inversely proportional to the wealth level. Specifically, the agent’s MV preference functional at time $t$ can be equivalently given by (3.4). It is easy to see that $g$ and $h$ are independent of $t$ and, for all $(v,x,y)\in[0,\infty)^{2}\times\mathbb{R}$,
@@
$-v_{a}(t)-y_{a}(t)$ $\leq-v_{a}(t)+\sqrt{v_{\lambda}(0)}\sqrt{v_{a}(t)}=\sqrt{v_{a}(t)}(\sqrt{v_{\lambda}(0)}-\sqrt{v_{a}(t)})\leq\frac{{v_{\lambda}(0)}}{4}.$
@@
$\tilde{a}(t)=\mathcal{P}_{t}\left(\tilde{h}(v_{\tilde{a}}(t),y_{\tilde{a}}(t))\right)=\mathcal{P}_{t}\left({h}(v_{\tilde{a}}(t),y_{\tilde{a}}(t))\right),\quad t\in(0,T).$
@@
Remark 4.4.
Consider the endogenous habit, scaled MV preferences in Section 4.2.2 of
Kryger, Nordfang, and Steffensen (2020), where they do not provide the proof of existence and uniqueness of solution of their ODE system. We define an equivalent preference functional $J(t,\pi)=\mathbb{E}_{t}\left[{W^{\pi}_{T}/W^{\pi}_{t}}\right]-\frac{\gamma}{2}\mathbb{E}_{t}\left[\left({W^{\pi}_{T}/W^{\pi}_{t}}-k(t)\right)^{2}\right]$. Suppose that $k(\cdot)$ is bounded and right-continuous. Using the same method as in Lemma 4.1 and Proposition 4.2 , we can show that there is a unique equilibrium $\bar{\pi}=(\sigma^{\top})^{-1}a$ in $\mathcal{D}$, where $a$ satisfies $a(t)=\left(\frac{1+\gamma k(t)}{\mathrm{e}^{y_{a}(t)+v_{a}(t)}}-1\right)\lambda(t)$, $t\in(0,T)$.
@@ 4.2 Equilibrium investment with random risk aversion
$h(x)={\mathbb{E}\left[\mathrm{e}^{-\frac{1}{2}\bm{R}x^{2}}\right]}\Big{/}{\mathbb{E}\left[\bm{R}\mathrm{e}^{-\frac{1}{2}\bm{R}x^{2}}\right]},\quad x\in[0,\infty).$ (4.3)
To study the Lipschitz continuity of $h$, we consider the following function: $f(x)=\mathbb{E}\left[\mathrm{e}^{-\frac{1}{2}\bm{R}x^{2}}\right]$, $x\in[0,\infty)$. The next lemma summarizes some well-known properties of $f$ and $h$; see, Klenke (2020, Example 6.29).
@@
If $\mathbb{E}\left[\bm{R}^{2}\right]<\infty$ and $h$ is bounded, then there is a unique equilibrium in $\mathcal{D}$.
@@
In the case when $\Gamma$ is a two-point distribution, Desmettre and Steffensen (2023) characterize the equilibrium by a three-dimensional ODE system666 In fact, with appropriate transformations, their ODE system can be converted into our integral equation. Specifically, there are typos in their solution $g^{\gamma}(t)$ and their equation (37). The solution $g^{\gamma}$ to their equation (35) should be $g^{\gamma}(t)=\exp\left(-\int_{t}^{T}\frac{1-\gamma}{\gamma}\left(-r+0.5\theta^{2}\frac{1}{(\varepsilon(s,\Gamma))^{2}}\gamma-\theta^{2}\frac{1}{\varepsilon(s,\Gamma)}\right)\mathrm{d}s\right)$ and hence their equation (37) for $\varepsilon(\cdot,\Gamma)$ should be $\varepsilon(t,\Gamma)=$ $\frac{\int\left(\exp\left(-\int_{t}^{T}\left(-r+0.5\theta^{2}\frac{1}{(\varepsilon(s,\Gamma))^{2}}\gamma-\theta^{2}\frac{1}{\varepsilon(s,\Gamma)}\right)\mathrm{d}s\right)\right)\gamma\mathrm{d}\Gamma(\gamma)}{\int\left(\exp\left(-\int_{t}^{T}\left(-r+0.5\theta^{2}\frac{1}{(\varepsilon(s,\Gamma))^{2}}\gamma-\theta^{2}\frac{1}{\varepsilon(s,\Gamma)}\right)\mathrm{d}s\right)\right)\mathrm{d}\Gamma(\gamma)}$ $=$ $\frac{\int\left(\exp\left(-\int_{t}^{T}\left(0.5\theta^{2}\frac{1}{(\varepsilon(s,\Gamma))^{2}}\gamma\right)\mathrm{d}s\right)\right)\gamma\mathrm{d}\Gamma(\gamma)}{\int\left(\exp\left(-\int_{t}^{T}\left(0.5\theta^{2}\frac{1}{(\varepsilon(s,\Gamma))^{2}}\gamma\right)\mathrm{d}s\right)\right)\mathrm{d}\Gamma(\gamma)}.$ (4.4) Consider the case $r=0$ and let $a(t)\triangleq\frac{\alpha}{\sigma}\frac{1}{\varepsilon(t,\Gamma)}=\frac{\theta}{\varepsilon(t,\Gamma)}$. Then Equation (4.4) is equivalent to our integral equation. without establishing the existence and uniqueness of the solution. Our approach, however, only requires that $\mathbb{E}\left[\bm{R}^{2}\right]<\infty$ and that $h$ is bounded to establish the existence and uniqueness of the solution. The boundedness of $h$ is fulfilled when the support of $\Gamma$ has a positive lower bound, that is, $\Gamma(\gamma_{0})=0$ for some $\gamma_{0}>0$.